# LLMTaxo: Leveraging Large Language Models for Constructing Taxonomy of Factual Claims from Social Media

### v1

@@ 5.2 Claim-Topic Evaluation
To assess how well the LLM-generated topics align with their corresponding factual claims, we performed evaluation on two aspects: accuracy and granularity. As we had already evaluated the taxonomy, which includes the relationship among broad, medium, and detailed topics in Section 5.1, we only evaluate claims and their leaf node topics. The detailed illustration of the evaluation metrics is:
@@
This criterion assesses how accurately the leaf node topics reflect the content and context of the corresponding factual claims. This involves determining if the topics are relevant and if they correctly represent the underlying information without misinterpretation or error.
@@
This criterion evaluates the specificity of the leaf node topics. This involves determining whether the topics are detailed enough to uniquely categorize and differentiate between factual claims, yet broad enough to maintain practical applicability across multiple claims.
After human evaluators rated each claim-topic pair, we calculated the average scores for each dataset separately for Zephyr and GPT-4o mini, as shown in Table 4. The evaluations highlight the distinct strengths of Zephyr and GPT-4o mini. Zephyr demonstrated consistent granularity across datasets but showed variable accuracy, with its strongest performance in the CV dataset. GPT-4o mini exhibited strong granularity, particularly in the CS dataset where it scored a high of 4.9. Although it generally scored lower than Zephyr’s in the CV dataset, it was competitive in the CC and CS datasets. The brevity and limited context of tweets in the CV dataset may have posed challenges for achieving high accuracy. In contrast, Facebook posts (CC and CS) likely provided more verbose and detailed content, which might have helped in achieving higher scores for GPT-4o mini.
@@ 5.3 Error Analysis
Based on the feedback from the human evaluators, errors in the taxonomies can be categorized into three types: 1) overlapping of topics, 2) lack of specificity in topic labels, and 3) generation of noisy data. For criteria receiving lower score in taxonomy evaluation, evaluators highlighted issues caused by overlapping and similar topic labels, such as “Vaccine Mandates” versus “COVID-19 Vaccine Mandates” or “Cybersecurity Levy on Transactions” versus “Cybersecurity Levy on Bank Transactions” Additionally, there are ambiguous labels, such as “Lawsuits” and “Cybersecurity,” which lack specificity. For the claim-topic pairs evaluation, some detailed topics directly replicate the factual claims. In addition, the LLMs may generate irrelevant data, producing detailed topics such as “not mentioned in the given post.” These points highlight areas where LLM performance can be refined to improve taxonomy accuracy and relevance.
@@ 6 Conclusion
In this study, we introduced $\mathsf{LLMTaxo}$, a novel framework that leverages LLMs to construct taxonomies of factual claims from social media. Through evaluations across three distinct datasets, our approach demonstrated effectiveness in organizing claims into hierarchical structures with broad, medium, and detailed topics. The results highlight the framework’s potential in reducing redundancy, improving information accessibility, and assisting researchers, fact-checkers, and policymakers in navigating online factual claims.
Limitation
Semantic variability presents another challenge, as similar claims may be expressed in different ways. Although the framework reduces the large portion of redundancy through clustering, subtle differences in wording can lead to semantically similar claims being categorized separately, potentially affecting the coherence of the generated taxonomy.
The framework’s reliance on LLMs such as Zephyr and GPT-4o mini also introduces potential limitations. These models, while powerful, may generate inconsistent or irrelevant topics for certain factual claims, which can impact the overall quality and reliability of the taxonomy. Despite efforts to automate the process, manual intervention remains necessary for refining and annotating learning examples. This introduces subjectivity, which may affect the consistency of the taxonomy.
Scalability poses an additional challenge. While $\mathsf{LLMTaxo}$ is designed to be adaptable across different domains, its performance may be constrained when applied to larger datasets with a broader range of topics. The computational demands associated with LLM-based topic generation could also limit its feasibility for large-scale applications.
@@ Appendix A Prompt for Topic Generation and Sample Results
There are $k$ learning examples used to guide the LLMs in generating a broad topic, a medium topic, and a detailed topic for each factual claim, as shown in Figure 2. Each prompt example contains a factual claim, a list of topic sets from the $k$ annotated factual claims (i.e., seed taxonomy), a question asking the LLMs to generate broad, medium, and detailed topics for the claim, and the answer to the question. In the question, the LLMs are instructed to prioritize generating topics from the existing topics. If none of the existing topics align well with the claim, the LLMs are then directed to generate new topics. This instruction ensures that the LLMs produce a limited number of topics. This prompt is iterated through all the factual claims to generate topics for them. A sample of the generated results from the three datasets are shown in Table 5.
@@
COVID-19 Vaccine John Stockton boldly suggests ’thousands’ of pro athletes died after Covid vaccine shot https://t.co/nXbt6Apm2q via @marca Vaccine Safety and Effectiveness Vaccine Side Effects Vaccine-Related Injuries and Deaths
@@
Appendix B Evaluation Metrics
@@
Appendix C GPT-4 Prompt for Evaluation
C.1 Prompt for Evaluating Taxonomy
I used LLMs to construct taxonomy and now I need to evaluate the taxonomy. I created some metrics to evaluate it. The [Taxonomy file name] uploaded contains the taxonomy with three-level topics. Please use the metrics in the [Metrics file name] to evaluate the taxonomy in [Taxonomy file name]. Please read each metric and understand them clearly, and then rate the metrics from 1-5, where 5 is the highest quality and 1 is the lowest. Please also provide judgments for your score and ignore the topic “Other” during evaluation.
C.2 Prompt for Evaluating Claim-Topic Pairs
<<EXAMPLES>>
<<EXAMPLE 1>>
Accuracy: 5. Granularity: 5.
<<EXAMPLE 2>>
Accuracy: 5. Granularity: 5.
<<EXAMPLE 3>>
Accuracy: 5. Granularity: 5.
<<EXAMPLE 4>>
Accuracy: 4. Granularity: 2.
<<EXAMPLE 5>>
Accuracy: 2. Granularity: 5.
<<END EXAMPLES>>
Claim: {claim}
Broad Topic: {broad_topic}
Medium Topic: {medium_topic}
Detailed Topic: {detailed_topic}

### v2

@@ 5.2 Claim-Topic Evaluation
To assess how well the factual claims are aligned with the topics assigned to them from the LLM-generated taxonomy, we performed evaluation on two aspects: accuracy and granularity. As Section 5.1 already evaluated the taxonomy itself, which includes the relationships among broad, medium, and detailed topics, we only evaluated claims and their leaf node topics. The detailed evaluation metrics are:
@@
This criterion assesses how accurately the leaf node topics reflect the content and context of the corresponding factual claims. This involves determining if the topics are relevant and if they correctly represent the underlying information in the claims without misinterpretation or error.
@@
This criterion evaluates the specificity of the leaf node topics. This involves determining whether the topics are detailed enough to uniquely categorize and differentiate between factual claims, yet broad enough to maintain practical applicability over multiple claims.
Nine human evaluators participated in the claim-topic evaluation. They were divided into three groups, each consisting of three individuals. Each group assessed the same set of 306 claim-topic pairs, comprising 34 pairs per model for each dataset. After completing the evaluations, we calculated the average scores for each model across the datasets, as presented in Table 4.
Overall, GPT-4o mini and Gemini demonstrated stronger performance in both accuracy and granularity compared to Zephyr, particularly on the CS dataset, where they achieved a granularity score of 4.9. According to the feedback from human evaluators, Zephyr often produced lengthy descriptions for detailed topics, occasionally repeating the content of the social media posts, which contributed to its lower granularity scores. However, Zephyr showed slightly higher accuracy in the CV dataset in the human evaluation, indicating it may perform better with shorter texts such as tweets. In contrast, Facebook posts (CC and CS) likely provided more verbose and detailed content, which may have benefited GPT-4o mini and Gemini.
To measure the inter-rater reliability, we computed Gwet’s AC2 scores for each group. The average scores of the three groups were 0.67 for accuracy and 0.75 for granularity.
@@ 5.3 Error Analysis
Based on the feedback from the human evaluators, errors in the taxonomies can be categorized into three types: 1) overlapping of topics, 2) lack of specificity in topic labels, and 3) generation of noisy data. For criteria receiving lower scores in taxonomy evaluation, evaluators highlighted issues caused by overlapping and similar topic labels, such as “Vaccine Mandates” versus “COVID-19 Vaccine Mandates” and “Cybersecurity Levy on Transactions” versus “Cybersecurity Levy on Bank Transactions.” Additionally, there are ambiguous labels, such as “Lawsuits” and “Cybersecurity,” which lack specificity. In the evaluation of claim-topic pairs, some detailed topics directly mirrored the factual claims. In addition, the LLMs may produce irrelevant outputs, such as detailed topics labeled “not mentioned in the given post.” These issues highlight areas where LLMs’ performance can be improved to enhance the accuracy and relevance of the taxonomy, as well as the necessity of post-processing the generated topics.
@@ 6 Conclusion
In this study, we introduced $\mathsf{LLMTaxo}$, a novel framework that leverages LLMs to construct taxonomies of factual claims from social media.Through evaluations across three distinct datasets, our approach demonstrated effectiveness in organizing claims into hierarchical structures with broad, medium, and detailed topics. The results highlight the framework’s potential in reducing redundancy, improving information accessibility, and assisting users such as researchers and fact-checkers in navigating online factual claims.
Limitations
The framework’s reliance on LLMs such as Zephyr and GPT-4o mini introduces potential limitations. These models, while powerful, may generate inconsistent or irrelevant topics for certain factual claims, which can impact the overall quality and reliability of the taxonomy. Despite efforts to automate the process, manual intervention remains necessary for refining and annotating learning examples. This introduces subjectivity, which may affect the consistency of the taxonomy.
Scalability poses an additional challenge. While $\mathsf{LLMTaxo}$ is designed to be adaptable for different topic domains, its performance may be constrained when applied to larger datasets with a broader range of topics. The computational demands associated with LLM-based topic generation could also limit its feasibility for large-scale applications.
@@
Acknowledgments
This work is partially supported by the United States National Science Foundation Award 2346261. We acknowledge the Texas Advanced Computing Center (TACC) for providing HPC resources that have contributed to the results reported in this paper. We are grateful to the human evaluators—Jacob Devasier, Abhishek Divakar Goudar, Juan Guajardo Gutierrez, Akshay Kumar Rayapet Madhusudhan, Alankrit Moses, Sai Sandeep Naraparaju, Jeremiah Pitts, Theodora Toutountzi, Mai Tran, Qing Wang, Devin Wingfield—for their thorough and dedicated work in taxonomy quality evaluation.
@@ Appendix A Prompt for Topic Generation and Sample Results
There are $k$ learning examples used to guide the LLMs in generating a broad topic, a medium topic, and a detailed topic for each factual claim, as shown in Figure 2. Each prompt example contains the list of topic tuples from the annotated factual claims in the $k$ learning examples (i.e., the seed taxonomy), a factual claim, a question asking the LLMs to generate broad, medium, and detailed topics for the claim, and the answer to the question. In the question, the LLMs are instructed to prioritize generating topics from the existing topics in the seed taxonomy. If none of the existing topics align well with the claim, the LLMs are then directed to generate new topics. This instruction ensures that the LLMs produce a limited number of topics. This prompt is iterated through all the factual claims to generate topics for them. A sample of the generated results from the three datasets (CV, CC, and CS) are shown in Table 5.
@@
COVID-19 Vaccine John Stockton boldly suggests ‘thousands’ of pro athletes died after Covid vaccine shot https://t.co/nXbt6Apm2q via @marca Vaccine Safety and Effectiveness Vaccine Side Effects Vaccine-Related Injuries and Deaths
@@
Appendix B GPT-4 Prompt for Evaluation
B.1 Prompt for Evaluating Taxonomy
I used LLMs to construct taxonomy and now I need to evaluate the taxonomy. I created some metrics to evaluate it. The [Taxonomy file name] uploaded contains the taxonomy with three-level topics. Please use the metrics in the [Metrics file name] to evaluate the taxonomy in [Taxonomy file name]. Please read each metric and understand them clearly, and then rate the metrics from 1-5, where 5 is the highest quality and 1 is the lowest. Please also provide justifications for your score and ignore the topic “Other” during evaluation.
B.2 Prompt for Evaluating Claim-Topic Pairs
<<EXAMPLES>>
<<EXAMPLE 1>>
Accuracy: 5. Granularity: 5.
<<EXAMPLE 2>>
Accuracy: 5. Granularity: 5.
<<EXAMPLE 3>>
Accuracy: 5. Granularity: 5.
<<EXAMPLE 4>>
Accuracy: 4. Granularity: 2.
<<EXAMPLE 5>>
Accuracy: 2. Granularity: 5.
<<END EXAMPLES>>
Claim: {claim}
Broad Topic: {broad_topic}
Medium Topic: {medium_topic}
Detailed Topic: {detailed_topic}
Appendix C Evaluation Metrics